OpenStudy (anonymous):

What is the midpoint of (–2, 0) and (4, –2)?

OpenStudy (anonymous):

midpoint formula is \[\frac{ x1+x2 }{ 2 };\frac{ y1=y2 }{ 2 }\]

OpenStudy (anonymous):

that should be y1+y2 my bad

OpenStudy (anonymous):

\[\frac{ x1+x2 }{ 2 },\frac{ y1+y2 }{2 }\]

OpenStudy (anonymous):

(-2,0 )=(x1,x2) (4,-2)=(y1,y2) substitute ...

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

hmm ...

OpenStudy (anonymous):

I got -1,1

OpenStudy (amistre64):

another option (just for show really :) ) is to simply move the points so one of them is at the origin (–2, 0) and (4, –2) +2 +2 ------- ------- (0,0) (6, -2) , half these amounts (3,-2) now move it back by subracting back out that 2 (3,-2) -2 ------ (1,-2)

OpenStudy (amistre64):

opps, some show, i forgot to half my -2 lol 1,-1 is fine

OpenStudy (anonymous):

Yeah you almost had me confused, thanks!

OpenStudy (anonymous):

ss ur ans is correct ,,,,

OpenStudy (anonymous):

thanks